OpenStudy (anonymous):

What is the equation of the line passing through the points (–2, 3) and (1, 4)?

OpenStudy (anonymous):

\[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] m is your slope \[(-2,3) ~ \implies ~ (x_1,y_1)\] \[(1,4) ~ \implies ~ (x_2, y_2)\]

OpenStudy (anonymous):

\[ \frac{y-(3)}{x-(-2)} = \frac{y-(4)}{x-(1)} \]If you solve for \(y\), you get the answer.

OpenStudy (anonymous):

This will actually work more easily:\[ \frac{y-(3)}{x-(-2)} = \frac{(3)-(4)}{(-2)-(1)} \]If you solve for \(y\), you get the answer.